ANALYSIS OF ROTATIONAL STRUCTURE IN THE HIGH-RESOLUTION INFRARED SPECTRUM OF CIS,CIS-1,4-DIFLUOROBUTADIENE

\underline{NORMAN C. CRAIG} and MICHAEL C. MOORE, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, OH 44074; TONY MASIELLO, Environmental and Molecular Sciences Laboratory, Pacific Northwest National Laboratory, P.O. Box 999, Mail Stop K8-88, Richland, WA 99352.

We seek an equilibrium structure cis,cis-1,4-difluorobutadiene. Rotational structure of a C-type band centered at 762.8 cm^{−1} in the high-resolution infrared spectrum (0.0015 cm^{−1}) has been analyzed as a first step. A sequence of strong hot bands of the torsional mode (78 cm^{−1}) complicate the analysis of this band. Provisional ground state rotational constants are reported. The spectrum of a second C-type band at 328 cm^{−1} may also be analyzable. Ground state rotational constants for a full set of isotopomers are needed. A procedure for synthesizing these species is being explored.